A 30 kg block is pushed up an inclined plane at an angle of 25° with a constant velocity of 4 m/s by a force of 280 N parallel to the plane.


What is the coefficient of kinetic friction between the block and the plane?



1
Expert's answer
2022-03-09T12:19:30-0500

Solution.

"m=30 kg;"

"\theta=25^o;"

"v=4 m\/s;"

"F=280N;"

"\mu-?;"

"F=F_f+mgsin\theta;"

"N=mgcos\theta;"

"F_f=\mu N=\mu mgcos\theta;"

"F=\mu mgcos\theta+mgsin\theta;"

"\mu mgcos\theta=F-mgsin\theta;"

"\mu=\dfrac{F-mgsin\theta}{mgcos\theta};"

"\mu=\dfrac{280-30\sdot9.81\sdot0.4226}{30\sdot9.81\sdot0.9063}=0.58;"


Answer:"\mu=0.58."